# A Geometric Unification of Generative AI with Manifold-Probabilistic Projection Models

### v1

@@ Abstract
The foundational premise of generative AI for images is the assumption that images are inherently low-dimensional objects embedded within a high-dimensional space. Additionally, it is often implicitly assumed that thematic image datasets form smooth or piecewise smooth manifolds. Common approaches overlook the geometric structure and focus solely on probabilistic methods, approximating the probability distribution through universal approximation techniques such as the kernel method. In some generative models, the low dimensional nature of the data manifest itself by the introduction of a lower dimensional latent space. Yet, the probability distribution in the latent or the manifold coordinate space is considered uninteresting and is predefined or considered uniform. This study unifies the geometric and probabilistic perspectives by providing a geometric framework and a kernel-based probabilistic method simultaneously. The resulting framework demystifies diffusion models by interpreting them as a projection mechanism onto the manifold of “good images”. This interpretation leads to the construction of a new deterministic model, the Manifold-Probabilistic Projection Model (MPPM), which operates in both the representation (pixel) space and the latent space. We demonstrate that the Latent MPPM (LMPPM) outperforms the Latent Diffusion Model (LDM) across various datasets, achieving superior results in terms of image restoration and generation.
@@ 1 Introduction
Restoration of images refers to the inverse process of generating a clean, meaningful, and non-corrupted image from a noisy, blurred, or otherwise degraded input. A critical aspect of this process involves the use of prior knowledge or a well-approximated distribution function over the set of clean images within a specific class. In this work, we propose the manifold assumption, which asserts that the set of desired images resides on a low-dimensional smooth manifold. We integrate this assumption with a probabilistic perspective. Specifically, we extend the conventional Monge patch description of the data manifold, typically provided by generative models such as autoencoders (AE) (ae_original, ), variational autoencoders (VAE) (kingma2022vae_1, ), and generative adversarial networks (GAN) (gan_original, ). Our approach augments this description by introducing a distance function that assigns, for each point in the pixel (ambient/representation) space, the distance to the closest point on the manifold. We treat here images as primary examples, but evidently it can be applied to any dataset that has this manifold structure. Next, we establish a connection between the geometric framework and the probabilistic perspective by introducing a geometric-based probability function and its kernel-based approximation. We further relate these approaches to diffusion-like methods, utilizing the score function to generate, in the ambient space, a vector field that directs each noisy or corrupted image towards the closest point on the manifold of clean images. By iteratively following this vector field, a diffusion-like flow is generated, guiding the corrupted image progressively towards a clean image residing on the manifold.
To accommodate the possibility of a nonuniform probability distribution on the manifold, we employ a kernel method that adjusts the diffusion-like flow to balance the trade-off between proximity to the manifold and the probability of a point on the manifold representing a clean and meaningful image. This integration of geometric principles with the kernel method constitutes the primary novelty of our approach. Furthermore, we extend these general concepts, the distance function, score, and diffusion-like flow, to operate within the latent space, thereby reducing computational complexity and enhancing the accuracy of the distance function. We evaluated our proposed method on the MNIST and SCUT-FBP5500 datasets, demonstrating superior performance compared to a leading method such as the Latent Diffusion Model (LDM) (rombach2022ldm, ).
@@ 1.1 Related Work
In recent years, the task of generating samples from a distribution that characterizes a specific dataset or target image has emerged as a critical challenge in machine learning. This problem has been extensively studied, with solutions primarily leveraging neural networks within deep learning frameworks. Many contemporary generative models operate under the implicit assumption that datasets comprise low-dimensional objects embedded within a high-dimensional space. However, the underlying geometry of the dataset is not always explicitly considered. For instance, variational autoencoders (VAEs) (kingma2022vae_1, ) and Generative Adversarial Networks (GANs) (gan_original, ) construct a functional mapping from the low-dimensional latent space to the high-dimensional pixel space. This functional mapping can be interpreted as a transformation from the manifold coordinate system to the pixel coordinate system. More recent approaches, such as diffusion models (diffusion_original, ; diffusion_2, ), adopt a more implicit perspective on manifold structure. Geometrically, these models can be viewed as learning a directional field that guides noisy points back to the data manifold, enabling iterative projection. The diffusion process gradually transforms random noise into realistic samples by iteratively denoising along paths that converge onto the data manifold.
@@
A central concept in many of these generative approaches is the Manifold Hypothesis (review_paper, ), which posits that real-world high-dimensional data, such as images, often concentrates near a low-dimensional manifold embedded within the ambient space. This geometric perspective provides a powerful conceptual framework for understanding generative models and has significantly influenced the design of numerous architectures and training objectives. Various other manifold-aware generative approaches have been proposed. Some methods explicitly model data as residing on specific manifolds. For instance, hyperspherical VAEs (davidson2022hypersphericalvae, ) and hyperbolic VAEs (mathieu2019, ) adapt generative models to handle data that naturally lies on non-Euclidean manifolds. Riemannian flow models (gemici2016normalizing, ; mathieu2020riemannian, ) incorporate Riemannian metrics into flow-based models to explicitly account for the intrinsic geometry of the data manifold. The relationship between manifold structure and probabilistic frameworks remains an active area of research. Normalizing flows (rezende2015variational, ) can be interpreted as learning diffeomorphisms between the data manifold and a simple base distribution. Score-based generative models (song2020, ) utilize the score function (the gradient of the log-density) to characterize the data distribution, establishing a direct connection to the geometry of the data manifold. Recent works on denoising diffusion models (ho2020denoising, ) can also be interpreted as learning a vector field that guides noisy samples back to the data manifold. Despite these advancements, there remains a gap in unifying the geometric and probabilistic perspectives in generative modeling.
@@ 2 Background and Theoretical Framework
2.1 Manifold Hypothesis
@@
In simple terms, the value at each pixel in the image (or in similar manifold-structured data) is a function of the $d$ parameters $z$.
2.2 Distance Function
@@
where for simplicity, we assume here that $\|\cdot\|$ denotes the Euclidean norm. In this high-dimensional representation space, the distance function provides a natural measure of the proximity of a point to the manifold. It is well known that $\mathcal{D}_{\mathcal{M}}$ satisfies the Eikonal equation $\nabla\mathcal{D}_{\mathcal{M}}(x)||=1$, with the natural boundary condition $\mathcal{D}_{\mathcal{M}}(x)=0$ for all $x\in\mathcal{M}$. Moreover, it is clear that $\nabla\mathcal{D}_{\mathcal{M}}(x)$ defines a vector field pointing in the direction of the shortest path to the manifold.
2.3 Image Generation/Denoising as Projection
This work unifies explicit and implicit manifold representations as the foundation of a generative image model. Specifically, we incorporate geometric constraints into a loss function that jointly trains a Monge patch Decoder/Generator $G$ and an extended encoder $F$ that operates throughout the ambient space - similar in spirit to a denoising autoencoder (DAE) denoising_original_vincent2008extracting . Although individual components of our algorithm are built on established architectures, the key contribution of our approach lies in the geometric perspective and the integration of the DAE framework with the learning of the distance function to the manifold. The three networks - Encoder $F$, Decoder/Generator $G$, and distance function $\mathcal{D}_{\mathcal{M}}$, are tightly coupled through the loss formulation, which is defined by the learned distance to the manifold.
2.4 Manifold Induced Probability on the Ambient Space
Building on this purely geometric consideration, we introduce a probabilistic model and demonstrate how a deterministic, stepwise diffusion-like model for projection onto the manifold is constructed using the score of the introduced probability function. Let us first assume that the probability of an arbitrary point in the ambient space being a clean image decreases exponentially with the distance from the manifold of clean and meaningful images. We adopt a simple model in which each image $x$ is assumed to be generated from the closest point $x^{*}$ on the manifold, with additive Gaussian noise, ${x}=x^{*}+\epsilon$, such that $\epsilon\sim\mathcal{N}(0,\sigma_{d}^{2})$. Therefore, the probability of x{x} to be on the manifold is given by
$P_{d}({x})=\frac{1}{Q_{d}}\exp\left({-\frac{\mathcal{D}^{2}_{\mathcal{M}}({x})}{2\sigma_{d}^{2}}}\right),$ (2)
where $Q_{d}$ is a normalization factor.
An alternative way to construct a probability distribution based on the manifold hypothesis is to describe each point x{x} in the ambient space as a noisy version of a point (or points) on the manifold, such that ${x}=G(z)+\epsilon$ is conditioned on $z$. The conditional probability is then given by
$P({x}|G(z))=\frac{1}{Q_{d}}\exp\left({-\frac{\|{x}-G(z)\|^{2}}{2\sigma_{d}^{2}}}\right).$ (3)
Based on the definition of the distance function, we derive the following expression using the maximum likelihood principle:
$P_{d}({x})=\max_{z}P({x}|G(z))=\frac{1}{Q_{d}}\exp\left({-\min_{z}\frac{\|{x}-G(z)\|^{2}}{2\sigma_{d}^{2}}}\right)=\frac{1}{Q_{d}}\exp\left(-\frac{\mathcal{D}^{2}_{\mathcal{M}}(x)}{2\sigma_{d}^{2}}\right).$ (4)
Note that $P_{d}$ implicitly assumes a uniform distribution of data points on the manifold, so the only factor that influences the (conditional) probability is the distance from the image $x$ in the ambient space to the manifold.
2.5 Kernel Methods
To account for a non-uniform probability distribution on the manifold, we define the probability function over the ambient space as a marginal distribution (see Appendix A for details):
$P_{\text{non-u}}(x)$ $=\int_{\mathbb{R}^{d}}P(x|G(z))P(z)dz,$
@@ 2.5 Kernel Methods
where the subscript "non-u" stands for non-uniform. In this formulation, the probability at $x$ is obtained by integrating contributions from all points on the manifold, where the conditional probability depends solely on the distance to the manifold and is thus purely geometric. Each contribution is weighted by $P(z)$, which represents the likelihood that the point $G(z)$ on the manifold corresponds to a clean image. Since the distribution $P(z)$ is unknown, we estimate it using a kernel density method. Specifically, we define:
[…]
@@
where $S$ is the set of latent code indices corresponding to clean images, and $Q_{\text{ker}}$ is the normalization constant. Note that $\sigma_{\text{ker}}$ is a hyperparameter that should be chosen carefully. In Fig. 2, we illustrate $P_{\text{ker}}(z)$. Clearly, the encoding of a generic image $x$ in the latent space, i.e., $F(x)$, may lie in a region with low probability. The probability of a point $x$ being an image depends on its distance to every point on the manifold, weighted by the probability of that point in the latent space. Using this kernel approximation together with the conditional probability from Eq. (3), we can thus approximate the probability function $P_{\text{non-u}}(x)$ as
[…]

### v2

@@ Abstract
Most models of generative AI for images assume that images are inherently low-dimensional objects embedded within a high-dimensional space. Additionally, it is often implicitly assumed that thematic image datasets form smooth or piecewise smooth manifolds. Common approaches overlook the geometric structure and focus solely on probabilistic methods, approximating the probability distribution through universal approximation techniques such as the kernel method. In some generative models the low dimensional nature of the data manifest itself by the introduction of a lower dimensional latent space. Yet, the probability distribution in the latent or the manifold’s coordinate space is considered uninteresting and is predefined or considered uniform. In this study, we address the problem of Blind Image Denoising (BID), and to some extent, the problem of generating images from noise by unifying geometric and probabilistic perspectives. We introduce a novel framework that improves upon existing probabilistic approaches by incorporating geometric assumptions that enable the effective use of kernel-based probabilistic methods. Furthermore, the proposed framework extends prior geometric approaches by combining explicit and implicit manifold descriptions through the introduction of a distance function. The resulting framework demystifies diffusion models by interpreting them as a projection mechanism onto the manifold of “good images”. This interpretation leads to the construction of a new deterministic model, the Manifold-Probabilistic Projection Model (MPPM), which operates in both the representation (pixel) space and the latent space. We demonstrate that the Latent MPPM (LMPPM) outperforms the Latent Diffusion Model (LDM) across various datasets, achieving superior results in terms of image restoration and generation.
@@ 1 Introduction
Restoration of images refers to the inverse process of generating a clean, meaningful, and non-corrupted image from a noisy, blurred, or other degraded input. A critical aspect of this process involves the use of prior knowledge or a well-approximated distribution function over the set of clean images within a specific class. In this work, we propose the manifold assumption, which asserts that the set of desired images resides on a low-dimensional smooth manifold. We integrate this assumption with a probabilistic perspective. Specifically, we extend the conventional Monge patch description of the data manifold, typically provided by generative models such as autoencoders (AE) (Rumelhart and McClelland, 1987), variational autoencoders (VAE) (Kingma and Welling, 2013), and generative adversarial networks (GAN) (Goodfellow et al., 2014). Our approach augments this description by introducing a distance function that assigns, for each point in the pixel (ambient/representation) space, the distance to the closest point on the manifold. We treat here images as primary examples, but evidently it can be applied to any dataset that has this manifold structure. Next, we establish a connection between the geometric framework and the probabilistic perspective by introducing a geometric-based probability function on the ambient space that is related naturally via the distance to points on the manifold to the probability distribution of data on the manifold. The latter is approximated on the latent space by kernel-based method and is pushed forward to the manifold via the generator. We further relate these approaches to diffusion-like methods, utilizing the score function to generate, in the ambient space, a vector field that directs each noisy or corrupted image towards a point on the manifold of clean images. The direction is a trade off between closeness to the corrupted image and points of higher probability on the manifold. By iteratively following this vector field, a diffusion-like flow is generated, guiding the corrupted image progressively towards a clean image residing on the manifold.
The introduction of the distance function that implicitly describe the manifold and the integration of geometric principles on the ambient space with the kernel method in the latent space constitute the primary novelties of our approach. Furthermore, we extend these general concepts, the distance function, score, and diffusion-like flow, to operate within the latent space, thereby reducing computational complexity and enhancing the accuracy of the distance function. We evaluated our proposed method on the MNIST, SCUT-FBP5500 and CelebA-HQ-256 datasets, demonstrating superior performance compared to a leading method such as the Latent Diffusion Model (LDM) (Rombach et al., 2022) and DiffBir (Lin et al., 2024).
@@ 1.1 Related Work
In recent years, the task of generating samples from a distribution that characterizes a specific dataset or target image has emerged as a critical challenge in machine learning. This problem has been extensively studied, with solutions primarily leveraging neural networks within deep learning frameworks. Many contemporary generative models operate under the implicit assumption that datasets comprise low-dimensional objects embedded within a high-dimensional space. However, the underlying geometry of the dataset is not always explicitly considered. For instance, autoencoders (AE) (Rumelhart and McClelland, 1987), variational autoencoders (VAEs) (Kingma and Welling, 2013) and Generative Adversarial Networks (GANs) (Goodfellow et al., 2014) construct a functional mapping from the low-dimensional latent space to the high-dimensional pixel space. This functional mapping can be interpreted as a transformation from the manifold coordinate system to the pixel coordinate system. More recent approaches, such as diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020a), adopt a more implicit perspective on manifold structure. Geometrically, these models can be viewed as learning a directional field that guides noisy points back to the data manifold, enabling iterative projection. The diffusion process gradually transforms random noise into realistic samples by iteratively denoising along paths that converge onto the data manifold.
@@
A central concept in many of these generative approaches is the Manifold Hypothesis (Loaiza-Ganem et al., 2024), which posits that real-world high-dimensional data, such as images, often concentrates near a low-dimensional manifold embedded within the ambient space. This geometric perspective provides a powerful conceptual framework for understanding generative models and has significantly influenced the design of numerous architectures and training objectives. Various other manifold-aware generative approaches have been proposed. Some methods explicitly model data as residing on specific manifolds. For instance, hyperspherical VAEs (Davidson et al., 2022) and hyperbolic VAEs (Mathieu et al., 2019) adapt generative models to handle data that naturally lies on non-Euclidean manifolds. Riemannian flow models (Gemici et al., 2016; Mathieu and Nickel, 2020) incorporate Riemannian metrics into flow-based models to explicitly account for the intrinsic geometry of the data manifold. Yet, in these flows the latent space is taken the size of the ambuent space ignoring the Manifold hypothesis that the dimension of the manifold is much smaller than the dimension of the ambient space. The relationship between manifold structure and probabilistic frameworks remains an active area of research. Normalizing flows (Rezende and Mohamed, 2015) can be interpreted as learning diffeomorphisms between the data manifold and a simple base distribution. Score-based generative models (Song and Ermon, 2020) utilize the score function (the gradient of the log-density) to characterize the data distribution, establishing a direct connection to the geometry of the data manifold. Recent works on denoising diffusion models (Ho et al., 2020b) can also be interpreted as learning a vector field that guides noisy samples back to the data manifold. Despite these advancements, there remains a gap in unifying the geometric and probabilistic perspectives in generative modeling. In particular the relations between the probability distribution in the ambient space, the probability distribution on the manifold and time/noise conditioning free diffusion like flow that is coherent with the above probability distributions and the structure of the manifold are still missing.
@@
1.2 Image Generation/Denoising as Projection
This work unifies explicit and implicit manifold representations as the foundation of a generative image model. Specifically, we incorporate geometric constraints into a loss function that jointly trains a Monge patch Decoder/Generator $G$ and an extended encoder $F$ that operates throughout the ambient space - similar in spirit to a denoising autoencoder (DAE) Vincent et al. (2008). Although individual components of our algorithm are built on established architectures, the key contribution of our approach lies in the geometric perspective and the integration of the DAE framework with the learning of the distance function to the manifold. The three networks - Encoder $F$, Decoder/Generator $G$, and distance function $\mathcal{D}_{\mathcal{M}}$, are tightly coupled through the loss formulation, which is defined by the learned distance to the manifold.
@@ 2 Background and Theoretical Framework
2.1 Geometry: Manifold Hypothesis - Explicit
@@
In simple terms, the value at each pixel in the image (or in similar manifold-structured data) is a function of the $d$ parameters $z$. In this work we use an autoencoder (or Denoising AE) where the encoder $F(x)$ maps an image $x\in\mathbb{R}^{D}$ to a point $z\in\mathbb{R}^{d}$. The generator $G$ is as above and is used as the map that defines the manifold.
Many works, in the context of deep learning, use this representation to analyze the data set as a Riemannian manifold. We will mention here, as examples, (Shao et al., 2018; Wang and Ponce, 2021) where geodesics and directions of meaningful changes on the manifold are studied. In (Chadebec and Allassonnière, 2022) the relation of the induced metric of the manifold was found to be approximated close enough to the encoded point of a clean image by the inverse covariance found in VAE.
2.2 Geometry: Distance Function - Implicit
@@
where $\|\cdot\|$ denotes the Euclidean norm. In this high-dimensional representation space, the distance function provides a natural measure of the proximity of a point to the manifold. It is well known that $\mathcal{D}_{\mathcal{M}}$ satisfies the Eikonal equation (Hamilton, 1828) $\nabla\mathcal{D}_{\mathcal{M}}(x)||=1$, with the natural boundary condition $\mathcal{D}_{\mathcal{M}}(x)=0$ for all $x\in\mathcal{M}$. Moreover, it is clear that $\nabla\mathcal{D}_{\mathcal{M}}(x)$ defines a vector field pointing in the direction of the shortest path to the manifold.
2.3 Probability: Ambient Space
On the probabilistic side, we follow works such as (Kadkhodaie et al., 2023) and (Sun et al., 2025). We begin by assuming a non-trivial distribution of clean data, denoted by $P_{c}(x)$, from which we have access to many samples namely, our dataset. We argue that for both Blind Image Denoising (BID) and image generation, the exact probability distribution is neither well-defined nor necessary. Indeed, we are not interested in the “true” probability of points outside the data manifold, i.e., degraded images. Rather, our goal is to construct a vector field that points from such degraded images toward the data manifold.
This objective can be achieved by defining a probability function that decreases monotonically with the distance from the manifold. More precisely, the probability assigned to a non-data point $x$ is defined so that the resulting score vector field points toward the data manifold and its more densely populated regions. To this end, we naturally define the conditional probability of a corrupted image $x$ given a clean image $x^{\prime}$ as
$P(x\mid x^{\prime})=f\bigl(d(x,x^{\prime})\bigr),$
where $d(x,x^{\prime})$ denotes the distance between the clean and corrupted images, and $f$ is a monotonically decreasing function.
For ease of analysis and computation, we choose $f$ to be Gaussian. This choice is not meant to suggest that the Gaussian is the “true” underlying function, but rather that it is sufficient for producing the desired vector field. As a second standard assumption, we take $d(x,x^{\prime})={\|x-x^{\prime}\|}$. These considerations lead to the following expression for the conditional probability:
$P_{\sigma}(x\mid x^{\prime})=\frac{1}{Q_{d}}\exp\left(-\frac{{\|x-x^{\prime}\|}^{2}}{2\sigma^{2}}\right),$ (2)
where $Q_{d}$ is the appropriate normalization constant.
The resulting probability distribution over the ambient space is then given by the well-known expression (Kadkhodaie et al., 2023; Sun et al., 2025):
$P(x;\sigma)=\int_{\mathbb{R}^{D}}P_{\sigma}(x\mid x^{\prime}),P_{c}(x^{\prime}),dx^{\prime}=\frac{1}{Q_{d}}\int_{\mathbb{R}^{D}}\exp\left(-\frac{{\|x-x^{\prime}\|}^{2}}{2\sigma^{2}}\right)P_{c}(x^{\prime})\,dx^{\prime},$ (3)
where $P_{c}(x^{\prime})$ denotes the probability density of clean images.
Although the Gaussian form may appear to impose a restriction to Gaussian noise, this formulation makes no explicit assumptions about the degradation process that maps $x^{\prime}$ to $x$. It merely assumes that the likelihood decreases exponentially with the distance between the corrupted and clean images. As a result, this approach yields a blind image denoising framework that does not require prior knowledge of the noise type or its magnitude. However, computing the score of this probability distribution is intractable in practice, as it requires evaluating an expectation over the entire ambient space.
2.4 Probability and Geometry: The Manifold Hypothesis
One of our main contributions is to introduce in this formulation the manifold hypothesis; namely, we propose that the probability of clean images is concentrated on a low-dimensional manifold
$P_{c}(x^{\prime})=\int P(G(z))\delta(x^{\prime}-G(z))dV_{\mathcal{M}},$ (4)
where $z=F(x^{\prime})$ is the latent representation of the clean image $x^{\prime}$. This means that $P_{c}(x^{\prime})=P(G(z))$ per manifold’s volume unit if $x^{\prime}\in\mathcal{M}$ and 0 otherwise. Substituting $P_{c}(x^{\prime})$ in Eq. (3) and after changing the order of integration, we obtain (the derivation is given in Appendix A.1):
${P_{\text{non-u}}(x;\sigma)}$ $=\int_{\mathbb{R}^{d}}P_{\sigma}(x\mid G(z))P(z)dz.$ (5)
Note that $P_{\text{non-u}}(x;\sigma)$ accounts for the nonuniform distribution of clean images on the manifold, and the integration is performed in the latent space. In the limit $\sigma\to 0$, only the point on the manifold closest to $x$ contributes significantly, and for small enough $\sigma$, denoted $\sigma_{d}$, we obtain:
$P_{\sigma_{d}}(x\mid G(z))\propto P_{d}({x}){:=}\frac{1}{Q_{d}}\exp\left({-\frac{\mathcal{D}^{2}_{\mathcal{M}}({x})}{2\sigma_{d}^{2}}}\right),$ (6)
where $Q_{d}$ is a normalization factor. In this limit, we obtain an “Energy-based model” where $E=\mathcal{D}_{\mathcal{M}}^{2}$. We can ignore the limiting process by which this EBM was derived and postulate it with a variance $\sigma_{d}^{2}$ of our choice. It is worth mentioning that learning directly the distance $\mathcal{D}_{\mathcal{M}}$ to the manifold makes the algorithm time/noise condition free. This distance encapsulates the noise/time approximation using the actual quantity of interest: the distance to the manifold of clean images. It therefore resolves another challenge associated with using diffusion models for image restoration (Sun et al., 2025).
2.5 Kernel Density Estimation in Latent Space
@@
In the formulation of $P_{\text{non-u}}(x;\sigma)$ derived above (5),(6), the probability at $x$ is obtained by integrating contributions from all points on the manifold, where the conditional probability depends solely on the distance to the manifold and is thus purely geometric. Each contribution is weighted by $P(z)$, which represents the likelihood that the point $G(z)$ on the manifold corresponds to a clean image.
Since the distribution $P(z)$ is unknown, we estimate it using a kernel density method (Rosenblatt, 1956; Parzen, 1962), a.k.a. ideal denoiser with delta mixture distribution / empirical distribution (Wang, 2024; Karras et al., 2020):
$P(z)\approx{P}_{\text{ker}}(z):=\frac{1}{Q_{\text{ker}}}\sum_{\alpha\in S}\exp\left({-\frac{\|z-z_{\alpha}\|^{2}}{2\sigma_{\text{ker}}^{2}}}\right),$ (7)
where $S$ is the set of latent code indices corresponding to clean images, and $Q_{\text{ker}}$ is the normalization constant. Note that $\sigma_{\text{ker}}$ is a hyperparameter that should be chosen carefully.
@@
In Fig. 2, we illustrate $P_{\text{ker}}(z)$. Clearly, the encoding of a generic image $x$ in the latent space, i.e., $F(x)$, may lie in a region with low probability. The probability of a point $x$ being an image depends on its distance to every point on the manifold, weighted by the probability of that point in the latent space. Using this kernel approximation together with the conditional probability (2), Eq. (5) takes the form
$P_{\text{non-u}}(x;{\sigma_{d}})\approx{\hat{P}_{\text{non-u}}(x;{\sigma_{d}}):=}\frac{1}{Q_{d}Q_{\text{ker}}}\sum_{\alpha\in S}\int_{\mathbb{R}^{d}}\exp\left({-\frac{\|{x}-G(z)\|^{2}}{2{\sigma_{d}}^{2}}}\right)\exp\left({-\frac{\|z-z_{\alpha}\|^{2}}{2\sigma_{\text{ker}}^{2}}}\right)dz.$ (8)